# Inner crust of neutron stars: Unified description of structure and superconductivity

### v1

@@ IV.2 Natural stress of the superconducting-superfluid crust in neutron stars
Let us generalize the expression for the force up to the case when the magnetic induction $\mathbf{B}$ has arbitrary angle $(\pi/2-\theta)$ with the neutron velocity perturbation. As a reminder, in Kobyakov2018 ) the angle was $\pi/2$, or equivalently $\theta=0$. The magnetic induction $\mathbf{B}$ is still restricted to the plane parallel to the slab surface and the momentum lag has the form
@@
where $n_{np0}$ is the entrainment number density inside the slab, $(\pi/2-\theta)$ is the angle between the magnetic field $\mathbf{B}_{0}$ and the vector of perturbation of the momentum lag $\delta\mathbf{w}$.
Since the slabs are independent ($\xi<d_{L}$), therefore the total force acting on a column of the 1N phase (as shown in Fig. 8 (a)) is given as a sum of contributions from each of the individual slabs, Eq. (207). It is convenient to express the momentum lag in terms of the velocity lag
$(\mathbf{\hat{y}}\cos\theta+\mathbf{\hat{x}}\sin\theta)\delta v=\mathbf{v_{p}}-\mathbf{v_{n}},$ (29)
@@
where $m_{0}^{*}$ is the so called isovector effective nucleon mass defined in Eq. (204) and
$\eta\equiv\frac{n}{n_{0}}.$ (32)
Combining Eqs. (28)-(31) and (207), we obtain
[…]
We assume the magnetic field (which in this case is almost the same as the magnetic induction) $B_{0}=5\times 10^{14}$ G and the velocity lag between neutrons and protons is $\delta v\sim 1\;{\rm cm\,s}^{-1}$, which corresponds to a rotational lag between superfluid neutrons and the normal part of the star of the order of $10^{-6}$ s-1. In this case, Eq. (33) can be written as an engineering formula:
@@
×$\times\left|1-\frac{m}{m_{0}^{*}}\right|u\frac{x\left(1-x\right)\eta^{2}}{1-2(1-\frac{m}{m_{0}^{*}})\eta}$
×$\times\left(\frac{D}{1\;{\rm cm}}\right)\left(\frac{B_{0}}{5\times 10^{14}\;{\rm G}}\right)\left(\frac{\delta v}{1\;{\rm cm\>s}^{-1}}\right).$
Noticing that at given $n_{b}$, $u$ is uniquely determined by the values of $n$ and $n_{no}$, and inserting the values for the variables $u$, $n$ and $x$ shown in Figs. 1-4, in Fig. 10 we plot the total force acting on a column of lasagna in the configuration shown in Figs. 8 (a), (c) for $\theta=\pi/2$, at $D=1\;{\rm cm}$, $B_{0}=5\times 10^{14}\;{\rm G}$, $\delta v=1\;{\rm cm\,s}^{-1}$, ${n_{0}}={0.1561\;{\rm fm}^{-3}}$ and $m_{0}^{*}/m=0.9466$.
Figure 10: Numerical evaluation of Eq. (34) in 1N phase with the variables $u$, $n$ and $x$ shown in Figs. 1-4 and with the other variables specified in the inset.
@@
It is worth to emphasize that we have assumed $D\sim 1\;{\rm cm}$ from the dimensional considerations and within the basic assumptions of the model, though likely $D\ll 1$ cm in the real neutron stars. As to the velocity perturbation $\delta v\sim 1\;{\rm cm\,s}^{-1}$, its magnitude has been assumed from the observational data on pulsars, which constrain the lag of the rotational frequencies of the superfluid and nonsuperfluid components to maximum values of the order of $10^{-6}\;{\rm s}^{-1}$ EysdenMelatos2010 .
V Summary and outlook
We have solved the complete set of coupled algebraic equations, Eqs. (4)–(7) to determine the equilibrium values, and thereby the equilibrium structure in LDM, of the number density of neutrons outside the nucleus or inside the bubble $n_{no}$, the number density of baryons $n$ and the proton fraction $x$ inside the nucleus or outside the bubble, and the nucleus radius $r_{N}$ or the bubble radius $r_{B}$, at given mean baryon density $n_{b}$. In this way we have calculated the energy spectrum of the hydrostatic configurations of our system that represents the neutron star matter in the inner crust.
[…]
We have found that at a fixed depth in the inner crust, which is characterized by a fixed average baryon number density $n_{b}$, there is a hydrostatic configuration with the minimum energy and, depending on the depth, there are other hydrostatic configurations with slightly higher energies. Our numerical calculations, based on a realistic nuclear interaction model, have shown that the energy difference is often smaller than the thermal energy per baryon. This finding suggests that in reality, matter in the inner crust at given depth might be polymorphic and contain coexisting various phases of the pasta structures (see Fig. 6 for more details).
The results of our initial estimates of the coherence length $\xi$, the slab width $2r_{N}$, the penetration depth of the magnetic field $\lambda$ and the Thomas-Fermi screening length of the electrons $k_{TFe}^{-1}$ suggest that the magnetic properties of the 1N phase in the pasta phase are nontrivial. As seen from Fig. 9, at $n_{b}\sim 0.48n_{0}$ there is an intersection of the line $\xi$ representing the coherence length and the line $d_{L}$ representing the distance between adjacent slabs. This implies that the characteristic radius of the normal core of quantized magnetic fluxtube for $0.37n_{0}<n_{b}<0.48n_{0}$ is smaller than the separation between adjacent slabs and therefore the magnetic flux is nonquantized, as shown in Fig. 8 (a). As a result, dynamical evolution of the magnetic flux in the material restricted to the geometry shown in Fig. 8 (a), is not accompanied by a dissipative dynamics of the nonuniform superconducting order parameter, in contrast to the configuration shown in Fig. 8 (b), where dynamical evolution of the magnetic flux is accompanied by the dissipative motion of the flux tube which intersects the slabs.
Our calculations reported in Fig. 5 for the workfunction to transfer a proton from the nucleus to the surrounding pure neutron matter, suggest that the energy landscape for the protons and their tunneling properties are more complicated than assumed in the earlier studies ZhangPethick2021 . Our calculations validate the approach within the anisotropic Ginzburg-Landau model used by Zhang and Pethick in ZhangPethick2021 for the mean baryon density range corresponding to $d_{L}/\xi<1$, which from Fig. 9 is at $0.48n_{0}<n_{b}<0.55n_{0}$. However, for $0.37n_{0}<n_{b}<0.48n_{0}$ the system should be described as a collection of individual two-dimensional slabs. In this case, the superconducting density is nonuniform and might be multiply-connected depending on the boundaries of the matter domains containing 1N pasta phase. Obviously, this physical situation is very different from a system where the superconducting matter is uniform and singly-connected as expected in the outer core of neutron stars.
As an astrophysical application, we have considered 1N phase in the magnetic field with a macroscopic flow of superfluid neutrons. In particular, we have studied the magnetohydrodynamic force arising from a perturbation of the neutron velocity $\delta v$ relative to the electron-proton plasma. In observable neutron stars with overlapping superfluid neutrons and protons in the interior, the superfluid lag $\delta v$ develops naturally, as a result of slowing down of the charged particles together with the pinned neutron vortices and pinned flux tubes, while keeping the superfluid neutron quantum rotational state unaltered during a significant period of time due to the neutron vortex pinning.
Noticing that the superconducting-superfluid entrainment energy contribution is a part of the magnetic piece of the free energy, we have revealed the connection between the stress tensor for the mixture with entrainment Sedrakian1995 ; Kobyakov2018 and the stress tensor previously considered for the case of uniform isotropic superconductors EassonPethick1977 . In this paper we have improved the numerical evaluation of the stress tensor in the inner crust of neutron stars, by using the parameters of the 1N phase calculated from microscopic physics and reported in Figs. 1-4.
With the new, realistic, estimates obtained in this work, we have corroborated the conclusion reached earlier in Kobyakov2018 : The characteristic superconductor force resulting from a specific combination of the pasta structure, the magnetic field and the neutron superfluid flow, with typical physical conditions expected in neutron stars, might acquire a magnitude that is large enough to shatter the crust. In this scenario, the rotational lag between the non-superfluid and superfluid components of the star, which is built up naturally in every observable neutron star with pinned superfluid vorticity, is a natural triggering agent for initiation of the crust shattering. The reasons that the force might acquire such a high magnitude are three-fold: (i) The non-quantized character of the magnetic flux in the configuration shown in Fig. 8 (a); (ii) a macroscopically large width $D$ of the 1N phase; (iii) the assumption that the slabs are ideally aligned in the azimuthal direction at any longitude with fixed latitude.
It is worth saying that there are directions for further improvement of our predictions regarding the inner crust structure. First, it would be interesting to include smaller corrections such as the neutrons adsorbed on the surface of nuclei and bubbles and some others. Second, the effects of a moderate strength magnetic field ($\sim 10^{15}$ Oe) on the inner crust structure should be studied, because the energy gaps in the hydrostatic energy spectrum have been found so small and might compare with the energy associated with the magnetic flux tube and the magnetized neutron vortex.
Qualitatively, in this paper we have confirmed by the numerical calculation that the electrical conductivity in the inner crust is not purely electronic. Since we have found that a significant part of the inner crust contains the 1N phase as the ground state, it is natural to expect that macroscopic domains of matter containing the 1-dimensional should be present in real neutron star crusts. Therefore, the electrical conductivity in some macroscopic parts of the inner crust in neutron stars is due to both the normal electrons and the superconducting protons.
@@ Acknowledgements
X.V. warmly acknowledges useful discussions with B. K. Sharma and M. Centelles. The work of D.K. reported in Secs. I-III was supported by Center of Excellence “Center of Photonics” sponsored by the Ministry of Science and Higher Education of Russia, contract № 075-15-2022-316. The work of D.K. reported in Sec. IV was supported by RSF project № 20-12-00268. X.V. acknowledges partial support from Grants No. PID2023-147112NB-C22 and No. CEX2019-000918-M (through the “Unit of Excellence María de Maeztu 2020-2023” award to ICCUB) from the Spanish MCIN/AEI/10.13039/501100011033.
@@ Appendix A The energy density functional and properties of uniform nuclear matter
In this work we use the Skyrme mean-field models constrained by the bulk matter equation of state computed using effective field theory and the ground-state energies of doubly magic nuclei introduced by Lim and Holt LimHolt2017 . In the ETF approximation the quantum kinetic energy density is replaced by the semiclassical one brack85 that at 2superscriptPlanck-constant-over-$\hbar^{2}$ order is expressed by a density functional that includes second order derivatives of the neutron proton densities. Within this approach the energy density for this kind of interactions reads:
@@
where $f_{q}$ ($q=n,p$) is the inverse of the nucleon effective mass for each nucleon species, which is defined as
@@
Here, $\tau_{q}$ and ${\bf J}_{q}$ are the 2limit-fromsuperscriptPlanck-constant-over-$\hbar^{2}-$order ETF kinetic energy density and the spin densities, respectively, which read
@@
In the LDM calculations the relevant part of the energy density functional (35) is the bulk part, which we write in a dimensionless form by scaling the energies and densities by the Fermi energy $\varepsilon_{0}$ and the saturation density, respectively. Mapping of the operator form of the interaction energy from Eqs. (3)-(4) of LimHolt2017 to the energy per baryon $\varepsilon$ in uniform nuclear matter can be done using the formula given by Eq. (1) of Dutra2012 , leading to:
@@
$n_{0}$ 0.1561141
$P_{\rm nuc}^{*}$ $3\times 10^{-6}$
$B$15.91
$K$ 239.3
$S$ 31.44
$L$ 42.07
[…]
@@
with $n_{t}$ given by Eq. (30) of LimHolt2017 .
@@
The nuclear binding energy per nucleon is given by
@@
Appendix B Nuclear clusters in LDM
The independent variables of the compressible liquid drop model for the nuclei are
$\{n_{b},\,n,\,x,\,r_{N},\,n_{no},\,n_{sk},\,n_{dc}\}.$ (47)
@@ Appendix B Nuclear clusters in LDM
$\{n,\,x,\,r_{N},\,n_{no},\,n_{sk},\,n_{dc}\}.$ (48)
The definition of the variables is given below.
@@ B.1 Energy density
where $r_{N}$ is the radius of the nucleus, $L\rightarrow\infty$ is the length of the rod-like nucleus and $L^{2}\rightarrow\infty$ is the area of the slab-like nucleus, see Fig. 11 (b). Here, $L\rightarrow\infty$ should be understood as large enough $L$ so that the Coulomb energy per unit length or area of nuclei and bubbles does not depend on $L$.
@@
where $r_{c}$ is the radius of the unit cell and $d$ is the dimensionality ($d=3$ for 3N, $d=2$ for 2N and $d=1$ for 1N). The volume of the unit cell is
$V_{c}=\left\{\begin{array}[]{c}({4}/{3})\pi r_{c}^{3},\quad(3N)\\ \pi r_{c}^{2}\times L,\quad(2N)\\ 2r_{c}\times L^{2},\quad(1N)\end{array}\right.$ (52)
Neutrons, however, are distributed within the entire cell: In the nucleus and in the volume of the cell excluding the volume of the nucleus. Moreover, there are neutrons adsorbed on the surface of the nucleus associated with the neutron skin and associated with the density curvature, i.e. the region where the second spatial derivative of the nucleon number density with respect to the coordinate normal to the nucleus surface is nonzero.
Thus, the total number of neutrons in the cell can be split into three parts:
@@
which corresponds to neutrons inside the nucleus, at the surface and outside the nucleus. We define the following neutron number densities.
@@
The surface neutrons (adsorbed on the surface of the nucleus) can be represented as neutrons associated with the neutron skin and with the surface width (the density curvature):
$N_{surf}=N_{sk}+N_{dc}.$ (58)
The cell-averaged number density of the skin neutrons is
$n_{sk}=\frac{N_{sk}}{V_{c}}.$ (59)
The cell-averaged number density of neutrons associated with the density curvature is
$n_{dc}=\frac{N_{dc}}{V_{c}}.$ (60)
Derivation of the governing variational equations of LDM assumes that $n_{sk}$ and $n_{dc}$ are independent variables. It is important that as soon as the variational equations are derived, the quantities $n_{sk}$ and $n_{dc}$ can be specified from the microscopic equations of the Skyrme model as following:
$n_{sk}=\left\{\begin{array}[]{c}{4\pi r_{N}^{2}s_{n}(n_{L}-n_{G})}{V_{c}^{-1}},\quad(3N)\\ {2\pi r_{N}s_{n}(n_{L}-n_{G})}{V_{c}^{-1}}\times L,\quad(2N)\\ {s_{n}(n_{L}-n_{G})}{V_{c}^{-1}}\times L^{2},\quad(1N)\end{array}\right.$ (61)
where $s_{n}=s_{n}(x)$ is the neutron skin thickness, and
$n_{dc}=\left\{\begin{array}[]{c}{8\pi r_{N}(b_{n}^{2}-b_{p}^{2}+s_{n}^{2})(n_{L}-n_{G})}{V_{c}^{-1}},\quad(3N)\\ {2\pi(b_{n}^{2}-b_{p}^{2}+s_{n}^{2})(n_{L}-n_{G})}{V_{c}^{-1}}\times L,\quad(2N)\\ 0,\quad(1N)\end{array}\right.$ (62)
where $b_{n}=b_{n}(x)$, $b_{p}=b_{p}(x)$ are the neutron surface width, proton surface width and $n_{L}=n_{L}(x)$ and $n_{G}=n_{G}(x)$ are some parameters; these quantities will be explicitly presented elsewhere.
$n_{b}=un+(1-u)n_{no}+n_{sk}+n_{dc}.$ (63)
@@
In LDM, the total energy density includes contributions to the energy per baryon, which we choose to express by the following 8 terms:
@@
$=w_{\mathrm{nuc}}+w_{\mathrm{p.surf}}+w_{\mathrm{C+L}}+w_{\mathrm{no}}+w_{\mathrm{e}}+w_{\mathrm{curv}}+w_{\mathrm{sk}}+w_{\mathrm{dc}}.$
where $w_{\mathrm{nuc}}$ is the bulk (uniform) energy density of the nuclear cluster. The terms $w_{\mathrm{p.surf}}$ and $w_{\mathrm{curv}}$ are the nuclear surface contributions (from the planar surface and from the curvature). The contribution from the dripped neutron liquid is given by $w_{\mathrm{no}}$ and the kinetic energy of the electron background by $w_{\mathrm{e}}$. The Coulomb energy term $w_{\mathrm{C+L}}$, collects the self-energy of protons in the nuclear cluster and of electrons distributed in the WS cell as well as the proton-electron lattice energy. Finally, the $w_{\mathrm{sk}}$ and $w_{\mathrm{dc}}$ provide the neutron energy due to neutrons accumulated at the surface of the nuclear cluster, being the former the contribution from the planar interface and the latter the correction due to its curvature.
The rest mass contribution due to baryons and the strong interaction contribution associated with the uniform matter inside the nucleus is given by
@@
where $m_{p}$ and $m_{n}$ are the proton and neutron rest masses, correspondingly, $c$ is the speed of light and $\varepsilon$ is the energy per baryon specified in the Appendix.
@@
where we have used the electrical neutrality condition in the unit cell, so the electron number density $n_{e}$ is given by
@@
The energy density associated with curvature of the thin surface of nuclear matter inside the nucleus for convenience can be called the curvature energy density $w_{\mathrm{curv}}$. This curvature energy density is proportional to both the surface area of the nucleus and to the mean principal curvature of the surface of the nucleus, which is $2/r_{N}$ for 3D clusters, $1/r_{N}$ for 2D clusters and 0 for 1D clusters. Thus, it has the form
@@
The energy density associated with the neutron skin can be written as
$w_{\mathrm{sk}}=n_{\mathrm{sk}}\mu_{\mathrm{sk}},$ (74)
where $\mu_{\mathrm{sk}}$ is defined by the variational equation with respect to $n_{\mathrm{sk}}$.
Analogously, the energy density associated with the nuclear density curvature can be written as
$w_{\mathrm{dc}}=n_{\mathrm{dc}}\mu_{\mathrm{dc}},$ (75)
where $\mu_{\mathrm{dc}}$ is defined by the variational equation with respect to $n_{\mathrm{dc}}$. This term is associated with the spatial curvature of the nuclear density related to the finite width of the nucleus surface.

### v2

@@ VII.2 Calculation of stress in a symmetric setting
Let us assume that the magnetic field $\mathbf{H}$ is pointed at an arbitrary angle $(\pi/2-\theta)$ relatively to the neutron velocity perturbation. Let us remind that in Kobyakov2018 ) the angle was $\pi/2$, or equivalently $\theta=0$. The magnetic induction $\mathbf{B}$ is still restricted to the plane parallel to the slab surface and the momentum lag has the form
@@
where $n_{np0}$ is the entrainment number density inside the slab, $(\pi/2-\theta)$ is the angle between the magnetic field $\mathbf{B}_{0}$ and the vector of perturbation of the momentum lag $\delta\mathbf{w}$, which points at the same direction as the neutron velocity perturbation.
It is convenient to express the momentum lag in terms of the velocity lag $\delta v$, defined as
$\mathbf{v_{p}}-\mathbf{v_{n}}=(\mathbf{\hat{y}}\cos\theta+\mathbf{\hat{x}}\sin\theta)\delta v,$ (42)
@@
where $\eta$ is given in Eq. (12). Combining Eqs. (41)-(44) and (39), we obtain
$\delta\mathbf{F}_{\rm tot}^{1\,{\rm cm}^{2}\times D}=-\mathbf{\hat{z}}\,2(1-\frac{m}{m_{0}^{*}})D\frac{e}{c}B_{0}n_{0}u\frac{x\left(1-x\right)\eta^{2}}{1-2(1-\frac{m}{m_{0}^{*}})\eta}\delta v.$ (45)
We assume the magnetic field $B_{0}=5\times 10^{14}$ G and the velocity lag between neutrons and protons $\delta v\sim 1\;{\rm cm\,s}^{-1}$. This value of $\delta v$ corresponds to a rotational lag between superfluid neutrons and the normal part of the star of the order of $10^{-6}$ s-1. In this case, Eq. (45) can be written as an engineering formula:
@@
$\times\left|1-\frac{m}{m_{0}^{*}}\right|u\frac{x\left(1-x\right)\eta^{2}}{1-2(1-\frac{m}{m_{0}^{*}})\eta}$
$\times\left(\frac{D}{1\;{\rm cm}}\right)\left(\frac{B_{0}}{5\times 10^{14}\;{\rm G}}\right)\left(\frac{\delta v}{1\;{\rm cm\>s}^{-1}}\right).$
Noticing that at given $n_{b}$, the variable $u$ is uniquely determined by the values of $n$ and $n_{no}$ given in Figs. 3 and 4, in Fig. 11 we plot the total force acting on a column of lasagna in the configuration shown in Figs. 8 (a), (c) for $\theta=\pi/2$, at $D=1\;{\rm cm}$, $B_{0}=5\times 10^{14}\;{\rm G}$, $\delta v=1\;{\rm cm\,s}^{-1}$, ${n_{0}}={0.1561\;{\rm fm}^{-3}}$ and $m_{0}^{*}/m=0.9466$.
Figure 11: Numerical evaluation of Eq. (46) in lasagna with the variables $u$, $n$ and $x$ shown in Figs. 1-4 and with the other parameters specified in the inset.
@@
It is worth to emphasize that we have assumed $D\sim 1\;{\rm cm}$ from the dimensional considerations and within the basic assumptions of the model, though likely $D\ll 1$ cm in real neutron stars. As to the velocity perturbation $\delta v\sim 1\;{\rm cm\,s}^{-1}$, its magnitude has been assumed from the observational data on pulsars, which constrain the lag of the rotational frequencies of the superfluid and nonsuperfluid components to maximum values of the order of $10^{-6}\;{\rm s}^{-1}$ EysdenMelatos2010 .
VIII Conclusions
We have solved the complete set of the variational equations of liquid drop model, Eqs. (4)–(7). We have computed self-consistently the surface tension of the interface that separates the nucleus from the pure neutron matter and the corresponding curvature correction, using the extended Thomas-Fermi approximation. In our calculations, the pairing energy gap has been extracted from the data obtained on the basis of same underlying nuclear interaction, the effective chiral field theory, as the one used in calculations of the bulk and the surface nuclear energies.
Our calculations show that inclusion of the curvature correction to the surface energy of nucleus changes the resulting energies of the pasta phases in such a way that the bubble phases completely disappear from the ground state, which is the state with the minimum internal energy per baryon among the entire set of pasta phases. We identify a low-energy subset of pasta phases, which includes the minimum-energy phase (shown by black solid line in Fig. 6) and the excited low-energy phases (shown by three types of color markers in Fig 6, corresponding to $T$ equal to $10^{8}$, $5\times 10^{8}$ and $10^{9}$ K). The excited phases have energy per baryon not exceeding that of the minimum-energy phase by more than $k_{B}T$.
We interpret the low-energy set as phases that appear during the crystallization process owing to the thermodynamic fluctuations. Thus, we conclude that the inner crust is likely to be polymorphic rather than a pure crystal. The spatial and temporal extension of the excited phases is to be calculated in the future. Appearance of the minimum-energy phases in the crust corresponds to the following pattern: 3N at lower mean baryon density $n_{b}$ followed by 2N, 1N and uniform at higher $n_{b}$. Lasagna is found in a significant part of the crust.
We have introduced a simple model of superconductivity in lasagna incorporated into liquid drop model. It follows from our numerical estimates that lasagna hosts a crossover between the discreet layered superconductivity and the anisotropic three-dimensional superconductivity. Based on our numerical results we have considered a symmetric stellar setting, as shown in Fig. 10, and have improved the earlier estimates of the stress tensor in lasagna, showing a possibility of shattering the crust by a combined effect of proton superconductivity, magnetic induction and the pulsar spin lag.
@@ Acknowledgements
X.V. warmly acknowledges useful discussions with B. K. Sharma and M. Centelles. X.V. acknowledges partial support from Grants No. PID2023-147112NB-C22 and No. CEX2019-000918-M (through the “Unit of Excellence María de Maeztu 2020-2023” award to ICCUB) from the Spanish MCIN/AEI/10.13039/501100011033. D.K. acknowledges partial support in the initial stage of the work from Center of Excellence “Center of Photonics” and RSF project 20-12-00268.
@@ Appendix A The energy density functional and properties of uniform nuclear matter
In this work we use the Skyrme mean-field model introduced by Lim and Holt LimHolt2017 , which is based on the effective field theory and constrained by the ground-state energies of doubly magic nuclei and observations of neutron stars. In the Skyrme mean-field approach, an extended Thomas-Fermi approximation is used, where the quantum kinetic energy density is replaced by the semiclassical one brack85 , which at $\hbar^{2}$ order is expressed by a density functional that includes second order derivatives of the neutron and proton densities. Within this approach the energy density for this kind of interactions reads:
@@
where $f_{q}$ ($q=n,p$) is the scaled inverse of the nucleon effective mass for each nucleon species, which is defined as
@@
Here, $\tau_{q}$ and ${\bf J}_{q}$ are the $\hbar^{2}$-order extended Thomas-Fermi kinetic energy density and the spin densities, respectively, which read
@@
In liquid drop model calculations the relevant part of the energy density functional (47) is the bulk part, which we write in a dimensionless form by scaling the energies and densities by the Fermi energy $\varepsilon_{0}$ and the saturation density, respectively. Mapping of the operator form of the interaction energy from Eqs. (3)-(4) of LimHolt2017 to the energy per baryon $\varepsilon$ in uniform nuclear matter can be done using the formula given in Eq. (1) of Dutra2012 , leading to:
@@
$n_{0}\;[{\rm fm}^{-3}]$ 0.1561
$P_{\rm nuc}\;[{\rm MeV\,fm}^{-3}]$ $3\times 10^{-6}$
$B\;[{\rm MeV}]$ -15.91
$K\;[{\rm MeV}]$ 239.3
$S\;[{\rm MeV}]$ 31.44
$L\;[{\rm MeV}]$ 42.07
Table 2: Nuclear interaction parameters are given by $t_{0,1,2,3,4}$, $x_{0,1,2,3,4}$, $\alpha_{1,2}$ LimHolt2017 . The empirical nuclear parameters $B$, $K$, $S$, $L$ are calculated from Eqs. (55)-(58). The nuclear symmetric saturation density $n_{0}$ is fine-tuned so that the empirical condition for the pressure, Eq. (54), is satisfied to a good precision.
@@
with $n_{t}$ given in Eq. (30) of LimHolt2017 .
@@
The nuclear binding energy per nucleon reads
@@
Appendix B Nucleus phases in liquid drop model
A complete set of variables of the compressible liquid drop model for the nucleus phases, which include 1N, 2N, 3N phases are
$\{n_{b},\,n,\,x,\,r_{N},\,n_{no}\}.$ (59)
@@ Appendix B Nucleus phases in liquid drop model
$\{n,\,x,\,r_{N},\,n_{no}\}.$ (60)
Their definition is given below.
@@ B.1 Energy density
where $r_{N}$ is the radius of the nucleus, $L\rightarrow\infty$ is the length of the rod-like nucleus and $L^{2}\rightarrow\infty$ is the area of the slab-like nucleus, see Fig. 11 (b). The limit $L\rightarrow\infty$ should be understood as large enough $L$ so that the Coulomb energy per unit length or area of nuclei and bubbles does not depend on $L$.
@@
where $r_{c}$ is the radius of the unit cell and $d=3$ for 3N, $d=2$ for 2N and $d=1$ for 1N.
Neutrons, however, are distributed within the entire cell, both inside and outside the nucleus. Moreover, there are neutrons adsorbed on the surface of the nucleus or the bubble. The adsorbed neutrons are associated with both the neutron skin and the density curvature, i.e. the region where the second spatial derivative of the nucleon number density with respect to the coordinate normal to the nucleus surface is nonzero. Thus, the total number of neutrons in the cell can be split into three parts:
@@
which corresponds to neutrons inside the nucleus, at the surface and outside the nucleus. In this paper, we neglect the neutron skin, $N_{surf}=0$.
@@
The cell-averaged number density of baryons can be written as
$n_{b}=un+(1-u)n_{no}.$ (69)
@@
In liquid drop model, the total energy density includes contributions to the energy per baryon, which we choose to express by the following 6 terms:
@@
$=w_{\mathrm{nuc}}+w_{\mathrm{p.surf}}+w_{\mathrm{C+L}}+w_{\mathrm{no}}+w_{\mathrm{e}}+w_{\mathrm{curv}}.$
where $w_{\mathrm{nuc}}$ is the bulk (uniform) energy density of the nuclear cluster. The terms $w_{\mathrm{p.surf}}$ and $w_{\mathrm{curv}}$ are the nuclear surface contributions, which represent a planar surface contribution and a curvature correction. The contribution from the dripped neutron liquid is given by $w_{\mathrm{no}}$ and the kinetic energy of the electron background by $w_{\mathrm{e}}$. The Coulomb energy term $w_{\mathrm{C+L}}$ collects the self-energy of protons in the nuclear cluster and of electrons distributed in the WS cell as well as the proton-electron lattice energy.
The rest mass contribution due to baryons and the strong interaction contribution associated with the uniform matter inside the nucleus reads
@@
where $m_{p}$ and $m_{n}$ are the proton and neutron rest masses, correspondingly, $c$ is the speed of light and $\varepsilon$ is the energy per baryon specified in Eq. (47).
@@
where we have used the electrical neutrality condition in the unit cell, so the electron number density $n_{e}$ reads
@@
The energy density associated with curvature of the thin surface of nuclear matter inside the nucleus is proportional to both the surface area of the nucleus and to the mean principal curvature of the surface of the nucleus, which is $2/r_{N}$ for nuclear clusters of 3N, $1/r_{N}$ for those of 2N and 0 for 1N. Thus, it has the form